A company purchased equipment for $325,000 on January 2, 2013. The

company expects the equipment to last for eight years or 60,000 hours

of operation, with an estimated salvage value of $25,000. During 2013,

the equipment was in operations for 8,000 hours, while in 2014 the

equipment was in operations for 8,700 hours. Compute the depreciation

expense relating to the equipment for 2013 and 2014 using the following

depreciation methods:

a. Straight-line

b. Double-declining-balance

c. Units-of-production.
Answer #1

Depreciation for year 2013 under straight line method

=( total cost - salvage value)/ estimated life

= (325000 - 25000)/8

= $ 37500

Depreciation for year 2014 under straight line method

= $ 37500

Depreciation for year 2013 under the written down value method

=cost ×25%

=325000×25%

= $ 81250

Depreciation for the year 2014 under the written down value method

= written down value × 25%

=243750 × 25%

= $ 60938

Depreciation for the year 2013 units of production

= no. Of hours x rate per activity

= 8000 × ((325000-25000)/60000)

=8000 × 5

= $ 40000

Depreciation for the year 2014 units of production

= no. Of hours × rate per activity

= 8700 × ((325000-25000)/60000)

= 8700×5

= $ 43500

Note

Rate of depreciation for WDV

= 100/life × 2

= 100/8 ×2

= 25 %

For any help please comment in the comment box.